Jaxson is a great free throw shooter. Over the last year he has made 88

%

88%

 of his free throws. If his probability stays consistent, what is the probability he will make his next three free throws? Show complete work on your worksheet!


1
Expert's answer
2021-02-24T06:53:12-0500

"n = 3 \\

p = 0.88 \\

(1 - p) = 0.12 \\

x = 3"

As per binomial distribution formula

"P(X = x) = nC^x \times p^x \times (1 - p)^{n - x}"

We need to calculate P(X = 3)

"P(X = 3) = 3C^3 \times 0.88^3 \times 0.12^0 \\

P(X = 3) = 0.6814"